[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Aug 10, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Aug 11 05:15:27 UTC 2017


HSU Shift Log for Fri Aug 11 2017 01:15:21 GMT-0400 (EDT)

Shift: R

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: The NTrans model loops or anomaly plots did not update for  
the 12 or 18Z cycle for any data source. These were Mike Brennan's old  
scripts. They are VERY helpful for operations!

AWIPS-2/NCP issues:

Other Techincal issues: ATT came out and checked the hurricane hotline.  
They found no issues locally and he "cleared" the lines and it looked fine  
to him. They closed the trouble ticket.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170811/4a3bb1fc/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list